OpenStudy (anonymous):

Where does the graph of y = 6x2 + 5x − 6 cross the x-axis? (−1/6, 0) and (6, 0) (1/6, 0) and (−6, 0) (2/3, 0) and (−, 0) (−2/3, 0) and (, 0)

OpenStudy (cwrw238):

you can try plugging in each of the above values to see if they fit or try to factor it

OpenStudy (cwrw238):

it will factor

OpenStudy (cwrw238):

to (3x - _ )(2x + _ ) can you fill in the blanks

OpenStudy (anonymous):

not sure how

OpenStudy (cwrw238):

you need 2 numbers a and b where a * b = -6 (the last term in trinomial) and where 3b - 2b = +5 ( the middle term +5x)

OpenStudy (cwrw238):

they are -2 and 3 so factors are (3x - 2)(2x + 3)

OpenStudy (cwrw238):

(3x - 2)(2x + 3) = 0 so solve for x